Eric is preparing to run in the Bay Marathon. One day, he ran and walked a total of 16 miles. If he ran one mile more than twice as far as he walked, how many miles did he run? Thank you for helping
Answer by checkley75(3666)

2x+1+X=16
3X=16-1
3X=15
X=15/3
X=5 ANSWER FOR THE DISTANCE HE WALKED.